OpenStudy (lgbasallote):

one of the problems in our very first integral calculus quiz..i never learned how to solve this..now might be a good time \[\large \int\limits_{}^{} \frac {sinx+cosx}{e ^{-x} + sinx} dx\]

OpenStudy (anonymous):

Igba.. come on u know the answer for sure!

OpenStudy (lgbasallote):

lol no =)) i hate integral calculus =))))

OpenStudy (anonymous):

forgive me for not knowing the integral thingy.. :)

OpenStudy (lgbasallote):

lol :))))) dont thingy me -__- :P i thought you took calculus already @dpaInc ?

OpenStudy (anonymous):

yep... just couldn't resist.

OpenStudy (accessdenied):

\[ \int \frac{sinx + cosx}{e^{-x} + sinx} dx \] \frac{numerator}{denominator}, maybe that should work

OpenStudy (lgbasallote):

ohhhh nice thanks!!! this works for everything right?? (btw..i probably wanted to know about the fraction thingy than the integral thingy haha)

OpenStudy (accessdenied):

\frac should work for all fractions, i think. it hasn't failed me yet. :D as for the actual integral, im not sure.

OpenStudy (lgbasallote):

is there a faq or discussion about the equation thingies @AccessDenied ? or maybe a way to know these secreat latex thingies or whatever they're called =))

OpenStudy (anonymous):